What is the potential gradient between two parallel plane conductors when their separation is 20 mm and a p.d. of 400 V is applied to them?

Calculate the force on an oil drop between the plates if the drop carries a charge of 8 × 10-19C.


1
Expert's answer
2021-01-10T18:30:53-0500

a) We can find the potential gradient between two parallel plane conductors from the formula:


"E=\dfrac{V}{d}=\dfrac{400\ V}{0.02\ m}=2.0\cdot10^4\ \dfrac{V}{m}."

b) We can find the force on an oil drop between the plates from the formula:


"F=qE=8.0\cdot10^{-19}\ C\cdot2.0\cdot10^4\ \dfrac{V}{m}=1.6\cdot10^{-14}\ N."

Answer:

a) "E=2.0\cdot10^4\ \dfrac{V}{m}."

b) "F=1.6\cdot10^{-14}\ N."